OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown?

OpenStudy (primeralph):

Question not complete.

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

..?

OpenStudy (primeralph):

Your question is missing some information.

OpenStudy (anonymous):

No it's not?

OpenStudy (anonymous):

28x- 12y= 16 and -6x+ 12y= -3 14x- 6y= 4 and -14x+ 28y= 1 -28x+ 12y= -16 and 28x- 56y= 14 those are the possible answers

OpenStudy (primeralph):

Yeah, needed to see those to answer it.

OpenStudy (primeralph):

Hold on.

OpenStudy (anonymous):

Okay.

OpenStudy (primeralph):

Is this B 14x- 6y= 4 and -14x+ 28y= 1 ?

OpenStudy (primeralph):

If the last digit is 1 and not 7, then that's your answer.

OpenStudy (anonymous):

Yes.

OpenStudy (primeralph):

B.

OpenStudy (anonymous):

Okay.